An 1200 kg purple pickup truck tows a small 300 kg trailer using a relatively lightweight tow rope as shown in the animation. The position of the trailer is given in 18.4m and time is in 16s.
Answer #1

Assuming that the force is constant, first of all you have to find the total displacement of the truck
S = x(final) - x(initial)

The displacement is connected to the acceleration of the system (a) and the time (t)

S = 1/2*a*t^2
a = 2S/(t^2) = 2*18.4/16^2 = 23/160 = 0.14375 m/s^2

The total force is

F = m(total)*a = (1200+300)a = 1500a = 1500*0.14375 = 215.625 N

Both the trailer and the truck move with the same acceleration.
The net force on the trailer is
F on trailer = m(trailer)*a = 300a = 43.125 N

and on the truck

F on truck = m(truck)*a = 1200a = 172.5 N